2#'s differ by 57, sum=185 the smallest out of the two numbers possibly could be 54,64,74: what would the equation be?
i tried 1x +2x+57=185 but i keep getting 42 and thats not one of the possible choices

Found 2 solutions by rfer, solver91311:
Answer by rfer(16322)

2x-57=185
2x=242
x=121
x-57=64
64+121=185

Answer by solver91311(24713)



Let represent the larger number and let represent the smaller number.

The difference is 57:



The sum is 185:



Add the two equations:








John